The sum of two number is 33. Thier different is 7. what are the two number
help please

Answer by stanbon(75887)

The sum of two number is 33.
Their difference is 7.
what are the two numbers?
====================
Equations:
x+y = 33
x-y = 7
---
Add and solve for "x":
2x = 40
x = 20
----
Substitute and solve for "y":
20 + y = 33
y = 13
-----------------
Cheers,
Stan H.